A hiker wants to determine the height of a tree. She measures the​ tree's shadow to be 17.1 meters long. Her own shadow at the same time is 1.8 meters long. If the hiker is 1.7 meters​ tall, how tall is the​ tree?

Answer by mananth(16949)


.
x/17.1 = 1.7/1.8
x = 17.1*1.7 /1.8
Height of tree = 16.15 m